OpenStudy (anonymous):

help me solve this with complete steps I'm very confused csch^-1 ((-1)/(2 sqrt(30)))

OpenStudy (anonymous):

i need to convert it into natural logarithums

OpenStudy (anonymous):

Refer to the attachment.

OpenStudy (anonymous):

sorry but i was looking for a detailed way to get to the answer

OpenStudy (dumbcow):

this should help.... http://math.info/image/192/inverse_hyperbolic_functions.gif \[csch^{-1} \frac{-1}{2 \sqrt{30}} = \sinh^{-1} (-2 \sqrt{30}) = \ln (-2\sqrt{30} +\sqrt{120+1})\]